# TalkSketch: Multimodal Generative AI for Real-time Sketch Ideation with Speech

### v1

@@ Introduction
Building on prior research in multimodal CUIs (hu2025gesprompt; cho2025persistent; masson2024directgpt), we expect that TalkSketch will help designers communicate intent more fluidly by combining verbal and visual cues. While existing conversational AI systems require users to articulate ideas through discrete text prompts, TalkSketch allows designers to “think aloud” while drawing, preserving the spontaneity of natural dialogue (he2025enhancingintentunderstandingambiguous; lan2025contextual; a18040214). This interaction form may reduce the cognitive effort involved in translating abstract ideas into prompts, a problem repeatedly identified in prompt-based creative tools (subramonyam2024bridging; wang2025enhancingcodellmsreinforcement). By aligning the AI’s interpretation with both speech and sketch input, users may experience the system as a more intuitive and responsive collaborator, one that feels closer to a design partner than a command-driven assistant. We therefore anticipate higher ratings of naturalness, communication clarity, and user control, reflecting an improved sense of mutual understanding between human and AI.
Supporting Creativity, Reflection, and Flow
@@
Consistent with prior work on creativity-support systems (peng2024designprompt; tao2025designweaver; lin2025inkspire; davis2025sketchai; kang2020towards; lan2025mappo), TalkSketch is designed to support early-stage ideation rather than polished production. We expect that integrating speech and sketch will enable designers to sustain creative flow by externalising thoughts continuously instead of interrupting the process to type or reformulate text. The AI Insight mechanism may further stimulate reflection by generating contextually relevant feedback from verbal and visual cues. Designers might pause to elaborate on their thinking, reinterpret sketches, or refine concepts based on AI suggestions, which echoes the dialogic creativity patterns observed in prior sketch-based systems (buxton2010sketching; suwa2022roles). This dynamic may lead to richer ideation traces and a stronger sense of co-evolution between human and AI ideas. Potential outcomes include higher perceived creativity support, increased exploratory behaviour, and a shift from command–response interaction to reflective conversation.
Conclusion and Future Work
@@
Future work will involve a controlled user study to systematically evaluate how TalkSketch affects the naturalness of human–AI interaction and the perceived creativity of the design process. This next step aims to provide empirical evidence for the benefits of sketch-and-speech interaction in multimodal generative design tools.

### v2

@@ 5.1 Enhancing Intent Expression and Interaction Naturalness
Building on prior research in multimodal CUIs [cho2025persistent, hu2025gesprompt, masson2024directgpt], we expect that TalkSketch will help designers communicate intent more fluidly by combining verbal and visual cues. While existing conversational AI systems require users to articulate ideas through discrete text prompts, TalkSketch allows designers to “think aloud” while drawing, preserving the spontaneity of natural dialogue [he2025enhancingintentunderstandingambiguous, lan2025contextual, a18040214]. This interaction form may reduce the cognitive effort involved in translating abstract ideas into prompts, a problem repeatedly identified in prompt-based creative tools [subramonyam2024bridging, wang2025enhancingcodellmsreinforcement]. By aligning the AI’s interpretation with both speech and sketch input, users may experience the system as a more intuitive and responsive collaborator, one that feels closer to a design partner than a command-driven assistant. We therefore anticipate higher ratings of naturalness, communication clarity, and user control, reflecting an improved sense of mutual understanding between human and AI.
5.2 Supporting Creativity, Reflection, and Flow
@@
Consistent with prior work on creativity-support systems [davis2025sketchai, kang2020towards, lan2025mappo, lin2025inkspire, peng2024designprompt, tao2025designweaver], TalkSketch is designed to support early-stage ideation rather than polished production. We expect that integrating speech and sketch will enable designers to sustain creative flow by externalising thoughts continuously instead of interrupting the process to type or reformulate text. The AI Insight mechanism may further stimulate reflection by generating contextually relevant feedback from verbal and visual cues. Designers might pause to elaborate on their thinking, reinterpret sketches, or refine concepts based on AI suggestions, which echoes the dialogic creativity patterns observed in prior sketch-based systems [buxton2010sketching, suwa2022roles]. This dynamic may lead to richer ideation traces and a stronger sense of co-evolution between human and AI ideas. Potential outcomes include higher perceived creativity support, increased exploratory behaviour, and a shift from command–response interaction to reflective conversation.
6 Limitations
This system has several limitations. First, because it depends on speech input, any transcription errors caused by background noise, unclear pronunciation, or the system mishearing words may lead to incorrect speech input. These mistakes may cause errors for the model to combine speech with sketches and to understand the user’s design intent. Second, in some cases, users may prefer not to speak when they are sketching alone. In such cases, the system may need to rely only on sketches and user prompting in the chatbot to support their needs, reducing a layer of speech information. However, as early-stage sketches are often incomplete or ambiguous, combining them with short or minimal prompts may make it difficult for the system to understand user needs compared to with speech information. Third, the system currently interprets speech and sketch inputs as a single combined chunk within a period of time, which inevitably leads to some information loss. Different parts of a spoken description may refer to different regions in the sketch, yet the system cannot yet distinguish these finer-grained correspondences. As a result, important links between what is said and what is drawn may be missed, limiting the precision with which user intent is understood.
7 Conclusion and Future Work
@@
Future work will involve a controlled user study to systematically evaluate how TalkSketch affects the naturalness of human–AI interaction and the perceived creativity of the design process. This next step aims to provide empirical evidence for the benefits of sketch-and-speech interaction in multimodal generative design tools. Beyond creativity support, it will also be valuable to explore how this interaction style can generalise to other settings, such as live demonstrations where rapid idea communication is essential, online classrooms where instructors sketch while explaining concepts, and collaborative design reviews where teams annotate evolving visuals. Examining these broader use cases may reveal additional opportunities for applying sketch-and-speech interaction as a more versatile interface paradigm.